OpenStudy (anonymous):

The life, X, of the StayBright light bulb is modelled by the probability density function\[f(x)=\left(\begin{matrix}2x^{-2} \\ 0\end{matrix}\right)\]\[x \ge 0\]otherwise. Where X is measures in thousands of hours Find the probability that a StayBrite bulb lasts longer than 1000 hours.

OpenStudy (anonymous):

f(x) is the function sine. Ok... How do you do that?

OpenStudy (anonymous):

sorry, it's 2e^-2

OpenStudy (anonymous):

so \[f(x)=2e ^{-2}\]

OpenStudy (anonymous):

and 0, yes

OpenStudy (anonymous):

But, it's basically just what you said.

OpenStudy (anonymous):

Oh no! Wait. It's e^-2x

OpenStudy (anonymous):

f(x)=2e^-2x

OpenStudy (anonymous):

Then integrate it from 1000 to infinty \[\int\limits_{1000}^{\infty} e ^{-2x}=-e ^{-2x}/2\]

OpenStudy (anonymous):

putting limits you get (e^-2000)/2.

OpenStudy (anonymous):

I thought it was -e?

OpenStudy (anonymous):

What exactly did you think was -e?

OpenStudy (anonymous):

No I mean, you wrote\[ -e^{-2x}\over2\]

OpenStudy (anonymous):

\[\int\limits_{}^{}e ^{ax} = e ^{ax}/a\]

OpenStudy (anonymous):

Yes, so, wouldn't it be -1/2 times e^-2x?

OpenStudy (anonymous):

thats exactly what i wrote.

OpenStudy (anonymous):

Yes, but when you wrote e^-2x putting in limits, would the e still carry the - sign?

OpenStudy (anonymous):

Substitue the limits with the minus sign you get the answer that i wrote only if a minus sign is there then we can reverse the limits and remove the minus sign(Property of definite integral.

OpenStudy (anonymous):

Ah Ok.

OpenStudy (anonymous):

So, how do you come to the answer?

OpenStudy (anonymous):

\[e ^{-\infty}=0\]

OpenStudy (anonymous):

I know the answer... Um, why do you do\[ e^{-\infty}=0?\]

OpenStudy (anonymous):

I just don't know how to get to the answer.

OpenStudy (anonymous):

After substitution, I got a strange answer...

OpenStudy (anonymous):

the limit is infinity right, so you get e^-2x= e power minus infinty which is 0.

OpenStudy (anonymous):

e^infinity= infinity so e^-infinity=1/infinity=0 (very vrey roughly). You can also observe from graph of e^-x